all numbers undefined for 2/4x+5
Divide and Simplify 5/t divided by 10/t
Subtract and simplify 13/d-8 - 13/8-d
add and simplify 7/15r + 4/9r

Answer by Alan3354(69443)

all numbers undefined for 2/4x+5
Is that a question?
-------------------
Divide and Simplify 5/t divided by 10/t
Multiply extremes and means - NUM = extremes = 5t
DEN = means = 10t
5t/10 = 1/2
-----------
Subtract and simplify 13/d-8 - 13/8-d
(d-8) = -(8-d)
--> 13/(d-8) + 13/(d-8)
= 26/(d-8)
Use parentheses. That could be (13/d) - 8 - (13/8) - d, an entirely different problem.
-----------------
add and simplify 7/15r + 4/9r
In class, they harp on the LCD for this type of problem, finding it, using it, etc.
You can multiply to the 2 DENs and use that, it will always work. You have to reduce the results, anyway.
Don't waste time on a test finding the LEAST common denominator, just find A common denominator (unless the problem asks for the LCD).
7/15r + 4/9r
Common DEN = 135r (no need to use r^2)
=63/135r + 60/135r
= 123/135r
= 41/45r